OpenStudy (nuccioreggie):

Which expression is equivalent to 73 ⋅ 7−5? 72 77 1 over 7 to the 2nd power 1 over 7 to the 7th power

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (tkhunny):

Show your work. Don't forget your Order of Operations.

OpenStudy (anonymous):

So \[\large7^3*\large7^{-5}\]?

OpenStudy (anonymous):

Hello? @nuccioreggie

OpenStudy (nuccioreggie):

Ok 1sec

OpenStudy (nuccioreggie):

Yea thats the question

OpenStudy (anonymous):

\[\Large7^3*\Large7^{-5}=\Large7^{3-5}\]

OpenStudy (nuccioreggie):

I think the answer is B

OpenStudy (anonymous):

And why do you think that?

OpenStudy (nuccioreggie):

Idk

OpenStudy (mathstudent55):

To multiply powers with the same base, give the same base and add the exponents. Examples: \(\large 5^8 \cdot 5^7 = 5^{8 + 7} = 5^{15} \) \(\large 3^{-5} \cdot 3^4 = 3^{-5 + 4} = 3^{-1} \)

OpenStudy (mathstudent55):

Also, after you multiply the powers together using the rule above, you need to know what to do with a negative exponent. For that use this rule: \(\large a^{-n} = \dfrac{1}{a^n} \) Example: \(\large 5^{-3} = \dfrac{1}{5^3} = \dfrac{1}{125} \)

OpenStudy (nuccioreggie):

Ok i just did that and i got the answer choice

OpenStudy (nuccioreggie):

A

OpenStudy (mathstudent55):

First, multiply the powers together. The base is 7. What do you get for the exponent?

OpenStudy (mathstudent55):

What is 3 + (-5) = ?

OpenStudy (nuccioreggie):

-2

OpenStudy (nuccioreggie):

Sorry if i typee the same thing twice its not moe its my computer

OpenStudy (mathstudent55):

Good. So far you are correct.